OpenStudy (anonymous):

"First make a substitution, then use integration by parts to evaluate the integral of 6sin(sqrt(x))dx." The only thing I can think of to sub is the sqrt(x) but then my dx is weird and I don't know where to go... help please?

OpenStudy (anonymous):

here is my guess, but i hate these \[u=\sqrt{x}\] \[du=\frac{dx}{2\sqrt{x}}\] \[2\sqrt{x}du=dx\]\[2udu=dx\]

OpenStudy (anonymous):

then integrate \[12\int u\sin(u)du\] by parts

OpenStudy (anonymous):

oh right you can sub the u back in for the sqrt(x), i remember doing it in class now, thank you so much

OpenStudy (anonymous):

the hint was "parts" otherwise i am sure i would not have guessed without it

OpenStudy (anonymous):

since \(x\sin(x)\) is a set up for parts yw